# Sample Compression Scheme Reductions

### v1

@@ Abstract
We present novel reductions from sample compression schemes in multiclass classification, regression, and adversarially robust learning settings to binary sample compression schemes. Assuming we have a compression scheme for binary classes of size $f(\mathrm{VC})$, where $\mathrm{VC}$ is the VC dimension, then we have the following results: (1) If the binary compression scheme is a majority vote or a stable compression scheme, then there exists a multiclass compression scheme of size $O\mathopen{}\left(f(\mathrm{d}_{\mathrm{G}})\right)$, where $\mathrm{d}_{\mathrm{G}}$ is the graph dimension. Moreover, for general binary compression schemes, we obtain a compression of size $O\mathopen{}\left(f(\mathrm{d}_{\mathrm{G}})\log|\mathcal{Y}|\right)$, where $\mathcal{Y}$ is the label space. (2) If the binary compression scheme is a majority vote or a stable compression scheme, then there exists an $\epsilon$-approximate compression scheme for regression over $[0,1]$-valued functions of size $O\mathopen{}\left(f(d_{\mathrm{P}})\right)$, where $d_{\mathrm{P}}$ is the pseudo-dimension. For general binary compression schemes, we obtain a compression of size $O\mathopen{}\left(f(d_{\mathrm{P}})\log(1/\epsilon)\right)$. These results would have significant implications if the sample compression conjecture, which posits that any binary concept class with a finite VC dimension admits a binary compression scheme of size $O(\mathrm{VC})$, is resolved [Littlestone and Warmuth, 1986, Floyd and Warmuth, 1995, Warmuth, 2003]. Our results would then extend the proof of the conjecture immediately to other settings. We establish similar results for adversarially robust learning and also provide an example of a concept class that is robustly learnable but has no bounded-size compression scheme, demonstrating that learnability is not equivalent to having a compression scheme independent of the sample size, unlike in binary classification, where compression of size $2^{O\mathopen{}\left(\mathrm{VC}\right)}$ is attainable [Moran and Yehudayoff, 2016].
@@ 1 Introduction
A common guiding principle in machine learning is to favor simpler hypotheses when possible, following Occam’s razor, which suggests that simpler models are more likely to generalize well. One approach to achieving simplicity, introduced by Littlestone and Warmuth [1986], Floyd and Warmuth [1995], is through a sample compression scheme for Probably Approximately Correct (PAC) learning [Valiant, 1984]. This framework simplifies the process of hypothesis learning by compressing the information needed to represent a learned model. This is done by encoding the hypothesis using a small subset of the original training data (along with a short bit string), known as the compression set, and a reconstruction function that recovers from this subset a sample-consistent hypothesis on the entire training set. The size of the compression set reflects the complexity of the learning task, with smaller sets implying simpler models (in some sense). A well-known example is the Support Vector Machine (SVM) algorithm, which constructs a halfspace in $\mathbb{R}^{d}$ using at most $d+1$ support vectors to represent its decision boundary. A significant open problem in binary classification, known as the sample compression conjecture, proposes that any concept class with a finite VC dimension admits a compression scheme of size $O(\mathrm{VC})$, where $\mathrm{VC}$ is the VC dimension [Warmuth, 2003]. A notable breakthrough by Moran and Yehudayoff [2016] demonstrated that every learnable binary concept class indeed admits a constant-size sample compression scheme (independent of the sample size), specifically of order $2^{O\mathopen{}\left(\mathrm{VC}\right)}$.
Beyond binary classification, sample compression schemes have also been explored in multiclass classification [Daniely et al., 2015, Daniely and Shalev-Shwartz, 2014, David et al., 2016, Brukhim et al., 2022, Pabbaraju, 2024]. In particular, it is known that multiclass learnability with a finite set of labels is equivalent to having a constant-size compression scheme, with a compression size of $2^{O\mathopen{}\left(\mathrm{d}_{\mathrm{G}}\right)}$ being achievable [David et al., 2016], where $\mathrm{d}_{\mathrm{G}}$ denotes the graph dimension of the concept class. However, more recently, Pabbaraju [2024] demonstrated that this equivalence no longer holds when the label set is infinite, and in such cases, any sample compression must grow at least logarithmically with the sample size.
@@
In this paper, we explore reductions from sample compression schemes in multiclass classification, regression, and adversarially robust learning settings to binary sample compression schemes. Assuming the existence of a compression scheme for binary classes of size $f(\mathrm{VC})$, we construct compression schemes of approximately the same order for these more general settings, where $\mathrm{VC}$ is replaced by the appropriate dimension, in both realizable and agnostic settings. Our results would have significant implications if the sample compression conjecture were resolved, as this would allow us to extend the proof of the conjecture to other settings immediately. We summarize our contributions as follows.
@@ 1.1 Our Results
Let $\mathcal{C}\subseteq\mathcal{Y}^{\mathcal{X}}$ be a multiclass concept class, with a finite graph dimension (denoted by $\mathrm{d}_{\mathrm{G}}$, see Definition 3.1). Note that any multiclass concept class (with finite label space) is learnable if and only if its graph dimension is finite.
  • We construct a sample compression scheme for multiclass classes of size $O\mathopen{}\left(f(\mathrm{d}_{\mathrm{G}}(\mathcal{C}))\log\mathopen{}\left|\mathcal{Y}\right|\right)$ (Theorem 3.2).
  • Assuming the reconstruction function of the compression scheme for binary classes either outputs a majority vote of concepts from $\mathcal{C}$ or selects a concept within the concept class $\mathcal{C}$ (proper compression), we construct a sample compression for $\mathcal{C}$ of size $O\mathopen{}\left(f(\mathrm{d}_{\mathrm{G}}(\mathcal{C}))\right)$, even when infinite label sets are allowed (Theorem 3.5). This result is interesting in light of the fact that a primary method for constructing binary schemes uses majority votes.
  • Assuming the existence of a stable sample compression scheme for binary concept classes, we construct a sample compression scheme for $\mathcal{C}$ of size $f(\mathrm{d}_{\mathrm{G}}(\mathcal{C}))$ for finite label sets (Theorem 3.7). Stable compression ensures that removing any point outside the compression set does not affect the output of the compression function. A notable example of such a compression scheme is the SVM algorithm. For infinite label sets, we introduce an infinitized version of compression schemes, which allows us to prove analogous results (Theorem 3.9 and Theorem 3.11). Surprisingly, we show that a finite VC dimension is not sufficient for such a compression to exist, unlike standard compression schemes. Additionally, we show that while finite Littlestone dimension implies infinitized compression, the converse does not hold.
@@
  • For bounded perturbation sets, letting $M=\sup_{x\in\mathcal{X}}\mathopen{}\left|\mathcal{U}(x)\right|$, we construct an adversarially robust sample compression of size $O\mathopen{}\left(f(\mathrm{VC}(\mathcal{C}))\log M\right)$ (Theorem 5.3). We get an improved compression size $O\mathopen{}\left(f\mathopen{}\left(\mathrm{VC}\right)\right)$ if $\mathcal{C}$ admits a binary stable compression scheme (Theorem 5.4).
@@ 2 Preliminaries
We say that $x_{1},\ldots,x_{n}\in\mathcal{X}$ are shattered by $\mathcal{C}\subseteq\mathopen{}\left\{0,1\right\}^{\mathcal{X}}$ if $\mathopen{}\left\{(c(x_{1}),\ldots,c(x_{n})):c\in\mathcal{C}\right\}=\{0,1\}^{n}$. The Vapnik-Chervonenkis (VC) dimension of a binary concept class $\mathcal{C}$, denoted by $\mathrm{VC}(\mathcal{C})$, is the largest nonnegative integer $n\in\mathbb{N}$ for which there exist $x_{1},\ldots,x_{n}\in\mathcal{X}\text{ that are shattered in }\mathcal{C}$.
The sample compression conjecture [Littlestone and Warmuth, 1986, Floyd and Warmuth, 1995, Warmuth, 2003] states that for classes with finite VC dimension, there exists a compression scheme of size $O(\mathrm{VC})$.
@@ 3 Compression for Multiclass Classification
In this section, we tackle the problem of multiclass compression by reducing it to the binary setting. In multiclass classification, the finiteness of the graph dimension of a concept class characterizes learnability when the label set is finite (albeit with nonoptimal sample complexity in general). David et al. [2016] demonstrated that, in such cases, a sample compression scheme of size $2^{O\mathopen{}\left(\mathrm{d}_{\mathrm{G}}\right)}$ can be constructed, with the compression size notably independent of the sample size. However, Pabbaraju [2024] showed that when the label set is infinite, there exist concept classes where any sample compression scheme must grow at least logarithmically with the sample size.
@@
The graph dimension of a multiclass concept class $\mathcal{C}$, denoted by $\mathrm{d}_{\mathrm{G}}(\mathcal{C})$, is the largest nonnegative integer $n\in\mathbb{N}$ for which there exist $x_{1},\ldots,x_{n}\in\mathcal{X}$ that are G-shattered by $\mathcal{C}$.
@@
that for binary concept classes with finite VC dimension $\mathrm{VC}<\infty$, there exists a sample compression scheme of size $f(\mathrm{VC})$. Then, for multiclass concept classes with a finite label set $|\mathcal{Y}|$ and a graph dimension $\mathrm{d}_{\mathrm{G}}<\infty$, there exists a sample compression scheme of size $O\mathopen{}\left(f(\mathrm{d}_{\mathrm{G}})\log|\mathcal{Y}|\right)$.
@@
Let $\mathcal{C}\subseteq\mathcal{Y}^{\mathcal{X}}$ be a multiclass concept class. Consider the inflated dataset $S_{\mathcal{Y}}$ (Equation 1) and the class $\mathcal{C}_{\mathcal{Y}}$ (Equation 2). Denote $\mathrm{VC}(\mathcal{C}_{\mathcal{Y}})=\mathrm{VC}$ and $\mathrm{d}_{\mathrm{G}}(\mathcal{C})=\mathrm{d}_{\mathrm{G}}$. It is straightforward to show that $\mathrm{VC}=\mathrm{d}_{\mathrm{G}}$. To show the $\leq$ direction, we have that for any $(x_{1},y_{1}),\ldots,(x_{n},y_{n})$ shattered by $\mathcal{C}_{\mathcal{Y}}$, all the $x_{i}$’s must be distinct (otherwise, there is an $(x,y_{j_{1}}),(x,y_{j_{2}})$ with $y_{j_{1}}\neq y_{j_{2}}$ that are shattered, and are both assigned value $1$ by a concept in $\mathcal{C}_{\mathcal{Y}}$, which implies that there exists an $c\in\mathcal{C}$ such that $c(x)=y_{j_{1}}$ and $c(x)=y_{j_{2}}$, which is not possible). Thus, $x_{1},\ldots,x_{n}$ are G-shattered in $\mathcal{C}$ via labels $y_{1},\ldots,y_{n}$. To show the $\geq$ direction, consider $x_{1},\ldots.,x_{n}$ G-shattered in $\mathcal{C}$ via labels $y_{1},\ldots,y_{n}$. It is clear that $(x_{1},y_{1}),\ldots,(x_{n},y_{n})$ are shattered in $\mathcal{C}_{\mathcal{Y}}$.
Suppose $\mathcal{C}_{\mathcal{Y}}$ has a binary compression scheme ($\kappa_{b}$,$\rho_{b}$), then we construct a compression scheme $(\kappa,\rho)$ for $\mathcal{C}$ as follows. Compression: Given a dataset $S=\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$ realizable by $\mathcal{C}$, construct $\kappa(S)$ as follows. Inflate $S$ to $\mathcal{S_{\mathcal{Y}}}$. Note that $S_{\mathcal{Y}}$ is realizable by $\mathcal{C}_{\mathcal{Y}}$, and $\mathrm{VC}=\mathrm{d}_{\mathrm{G}}$. We can apply $\kappa_{b}$ to $S_{\mathcal{Y}}$ to get a compression of size $f(\mathrm{d}_{\mathrm{G}})$. The compression points will be of the form $((x,y),z),z\in\{0,1\}$. For the points where $z=1$, we have that $(x,z)=(x_{i},y_{i})$ for some $i$, so we can add that to $\kappa(S)$, contributing $1$ for each. For the points where $z=0$, we need $\log|\mathcal{Y}|$ bits to add that to $\kappa(S)$. Thus, our compression size will be $\leq f(\mathrm{d}_{\mathrm{G}})+f(\mathrm{d}_{\mathrm{G}})\log|\mathcal{Y}|=O(f(\mathrm{d}_{\mathrm{G}})\log|\mathcal{Y}|)$. Reconstruction: Our compression has enough information for us to retrieve the result of $\kappa_{b}(\mathcal{C}_{\mathcal{Y}})$. We can directly apply $\rho_{b}$ on this to get the desired result.
@@
Suppose all binary concept classes with VC dimension $\mathrm{VC}$ have a sample compression scheme of size $f(\mathrm{VC})$. Does every multiclass class with graph dimension $\mathrm{d}_{\mathrm{G}}$ have a sample compression scheme of size $O\mathopen{}\left(f(\mathrm{d}_{\mathrm{G}})\right)$?

### v2

@@ Abstract
We present novel reductions from sample compression schemes in multiclass classification, regression, and adversarially robust learning settings to binary sample compression schemes. Assuming we have a compression scheme for binary classes of size $f(d_{\mathrm{VC}})$, where $d_{\mathrm{VC}}$ is the VC dimension, then we have the following results: (1) If the binary compression scheme is a majority vote or a stable compression scheme, then there exists a multiclass compression scheme of size $O\mathopen{}\left(f(d_{\mathrm{G}})\right)$, where $d_{\mathrm{G}}$ is the graph dimension. Moreover, for general binary compression schemes, we obtain a compression of size $O\mathopen{}\left(f(d_{\mathrm{G}})\log|\mathcal{Y}|\right)$, where $\mathcal{Y}$ is the label space. (2) If the binary compression scheme is a majority vote or a stable compression scheme, then there exists an $\epsilon$-approximate compression scheme for regression over $[0,1]$-valued functions of size $O\mathopen{}\left(f(d_{\mathrm{P}})\right)$, where $d_{\mathrm{P}}$ is the pseudo-dimension. For general binary compression schemes, we obtain a compression of size $O\mathopen{}\left(f(d_{\mathrm{P}})\log(1/\epsilon)\right)$. These results would have significant implications if the sample compression conjecture, which posits that any binary concept class with a finite VC dimension admits a binary compression scheme of size $O(d_{\mathrm{VC}})$, is resolved [Littlestone and Warmuth, 1986, Floyd and Warmuth, 1995, Warmuth, 2003]. Our results would then extend the proof of the conjecture immediately to other settings. We establish similar results for adversarially robust learning and also provide an example of a concept class that is robustly learnable but has no bounded-size compression scheme, demonstrating that learnability is not equivalent to having a compression scheme independent of the sample size, unlike in binary classification, where compression of size $2^{O\mathopen{}\left(d_{\mathrm{VC}}\right)}$ is attainable [Moran and Yehudayoff, 2016].
@@ 1 Introduction
A common guiding principle in machine learning is to favor simpler hypotheses when possible, following Occam’s razor, which suggests that simpler models are more likely to generalize well. One approach to achieving simplicity, introduced by Littlestone and Warmuth [1986], Floyd and Warmuth [1995], is through a sample compression scheme for Probably Approximately Correct (PAC) learning [Valiant, 1984]. This framework simplifies the process of hypothesis learning by compressing the information needed to represent a learned model. This is done by encoding the hypothesis using a small subset of the original training data (along with a short bit string), known as the compression set, and a reconstruction function that recovers from this subset a sample-consistent hypothesis on the entire training set. The size of the compression set reflects the complexity of the learning task, with smaller sets implying simpler models (in some sense). A well-known example is the Support Vector Machine (SVM) algorithm, which constructs a halfspace in $\mathbb{R}^{d}$ using at most $d+1$ support vectors to represent its decision boundary. A significant open problem in binary classification, known as the sample compression conjecture, proposes that any concept class with a finite VC dimension admits a compression scheme of size $O(d_{\mathrm{VC}})$, where $d_{\mathrm{VC}}$ is the VC dimension [Warmuth, 2003]. A notable breakthrough by Moran and Yehudayoff [2016] demonstrated that every learnable binary concept class indeed admits a constant-size sample compression scheme (independent of the sample size), specifically of order $2^{O\mathopen{}\left(d_{\mathrm{VC}}\right)}$.
Beyond binary classification, sample compression schemes have also been explored in multiclass classification [Daniely et al., 2015, Daniely and Shalev-Shwartz, 2014, David et al., 2016, Brukhim et al., 2022, Pabbaraju, 2024]. In particular, it is known that multiclass learnability with a finite set of labels is equivalent to having a constant-size compression scheme, with a compression size of $2^{O\mathopen{}\left(d_{\mathrm{G}}\right)}$ being achievable [David et al., 2016], where $d_{\mathrm{G}}$ denotes the graph dimension of the concept class. However, more recently, Pabbaraju [2024] demonstrated that this equivalence no longer holds when the label set is infinite, and in such cases, any sample compression must grow at least logarithmically with the sample size.
@@
In this paper, we explore reductions from sample compression schemes in multiclass classification, regression, and adversarially robust learning settings to binary sample compression schemes. Assuming the existence of a compression scheme for binary classes of size $f(d_{\mathrm{VC}})$, we construct compression schemes of approximately the same order for these more general settings, where $d_{\mathrm{VC}}$ is replaced by the appropriate dimension, in both realizable and agnostic settings. Our results would have significant implications if the sample compression conjecture were resolved, as this would allow us to extend the proof of the conjecture to other settings immediately. We summarize our contributions as follows.
@@ 1.1 Our Results
Let $\mathcal{C}\subseteq\mathcal{Y}^{\mathcal{X}}$ be a multiclass concept class, with a finite graph dimension (denoted by $d_{\mathrm{G}}$, see Definition 3.1). Note that any multiclass concept class (with finite label space) is learnable if and only if its graph dimension is finite.
  • We construct a sample compression scheme for multiclass classes of size $O\mathopen{}\left(f(d_{\mathrm{G}}(\mathcal{C}))\log\mathopen{}\left|\mathcal{Y}\right|\right)$ (Theorem 3.2).
  • Assuming the reconstruction function of the compression scheme for binary classes either outputs a majority vote of concepts from $\mathcal{C}$ or selects a concept within the concept class $\mathcal{C}$ (proper compression), we construct a sample compression for $\mathcal{C}$ of size $O\mathopen{}\left(f(d_{\mathrm{G}}(\mathcal{C}))\right)$, even when infinite label sets are allowed (Theorem 3.5). This result is interesting in light of the fact that a primary method for constructing binary schemes uses majority votes.
  • Assuming the existence of a stable sample compression scheme for binary concept classes, we construct a sample compression scheme for $\mathcal{C}$ of size $f(d_{\mathrm{G}}(\mathcal{C}))$ for finite label sets (Theorem 3.7). Stable compression ensures that removing any point outside the compression set does not affect the output of the compression function. A notable example of such a compression scheme is the SVM algorithm. For infinite label sets, we introduce an infinitized version of compression schemes, which allows us to prove analogous results (Theorem 3.9 and Theorem 3.11). Surprisingly, we show that a finite VC dimension is not sufficient for such a compression to exist, unlike standard compression schemes. Additionally, we show that while finite Littlestone dimension implies infinitized compression, the converse does not hold.
@@
  • For bounded perturbation sets, letting $M=\sup_{x\in\mathcal{X}}\mathopen{}\left|\mathcal{U}(x)\right|$, we construct an adversarially robust sample compression of size $O\mathopen{}\left(f(d_{\mathrm{VC}}(\mathcal{C}))\log M\right)$ (Theorem 5.3). We get an improved compression size $O\mathopen{}\left(f\mathopen{}\left(d_{\mathrm{VC}}\right)\right)$ if $\mathcal{C}$ admits a binary stable compression scheme (Theorem 5.4).
@@ 2 Preliminaries
We say that $x_{1},\ldots,x_{n}\in\mathcal{X}$ are shattered by $\mathcal{C}\subseteq\mathopen{}\left\{0,1\right\}^{\mathcal{X}}$ if $\mathopen{}\left\{(c(x_{1}),\ldots,c(x_{n})):c\in\mathcal{C}\right\}=\{0,1\}^{n}$. The Vapnik-Chervonenkis (VC) dimension of a binary concept class $\mathcal{C}$, denoted by $d_{\mathrm{VC}}(\mathcal{C})$, is the largest nonnegative integer $n\in\mathbb{N}$ for which there exist $x_{1},\ldots,x_{n}\in\mathcal{X}\text{ that are shattered in }\mathcal{C}$.
The sample compression conjecture [Littlestone and Warmuth, 1986, Floyd and Warmuth, 1995, Warmuth, 2003] states that for classes with finite VC dimension, there exists a compression scheme of size $O(d_{\mathrm{VC}})$.
@@ 3 Compression for Multiclass Classification
In this section, we tackle the problem of multiclass compression by reducing it to the binary setting. In multiclass classification, the finiteness of the graph dimension of a concept class characterizes learnability when the label set is finite (albeit with nonoptimal sample complexity in general). David et al. [2016] demonstrated that, in such cases, a sample compression scheme of size $2^{O\mathopen{}\left(d_{\mathrm{G}}\right)}$ can be constructed, with the compression size notably independent of the sample size. However, Pabbaraju [2024] showed that when the label set is infinite, there exist concept classes where any sample compression scheme must grow at least logarithmically with the sample size.
@@
The graph dimension of a multiclass concept class $\mathcal{C}$, denoted by $d_{\mathrm{G}}(\mathcal{C})$, is the largest nonnegative integer $n\in\mathbb{N}$ for which there exist $x_{1},\ldots,x_{n}\in\mathcal{X}$ that are G-shattered by $\mathcal{C}$.
@@
that for binary concept classes with finite VC dimension $d_{\mathrm{VC}}<\infty$, there exists a sample compression scheme of size $f(d_{\mathrm{VC}})$. Then, for multiclass concept classes with a finite label set $|\mathcal{Y}|$ and a graph dimension $d_{\mathrm{G}}<\infty$, there exists a sample compression scheme of size $O\mathopen{}\left(f(d_{\mathrm{G}})\log|\mathcal{Y}|\right)$.
@@
Let $\mathcal{C}\subseteq\mathcal{Y}^{\mathcal{X}}$ be a multiclass concept class. Consider the inflated dataset $S_{\mathcal{Y}}$ (Equation 1) and the class $\mathcal{C}_{\mathcal{Y}}$ (Equation 2). Denote $d_{\mathrm{VC}}(\mathcal{C}_{\mathcal{Y}})=d_{\mathrm{VC}}$ and $d_{\mathrm{G}}(\mathcal{C})=d_{\mathrm{G}}$. It is straightforward to show that $d_{\mathrm{VC}}=d_{\mathrm{G}}$. To show the $\leq$ direction, we have that for any $(x_{1},y_{1}),\ldots,(x_{n},y_{n})$ shattered by $\mathcal{C}_{\mathcal{Y}}$, all the $x_{i}$’s must be distinct (otherwise, there is an $(x,y_{j_{1}}),(x,y_{j_{2}})$ with $y_{j_{1}}\neq y_{j_{2}}$ that are shattered, and are both assigned value $1$ by a concept in $\mathcal{C}_{\mathcal{Y}}$, which implies that there exists an $c\in\mathcal{C}$ such that $c(x)=y_{j_{1}}$ and $c(x)=y_{j_{2}}$, which is not possible). Thus, $x_{1},\ldots,x_{n}$ are G-shattered in $\mathcal{C}$ via labels $y_{1},\ldots,y_{n}$. To show the $\geq$ direction, consider $x_{1},\ldots.,x_{n}$ G-shattered in $\mathcal{C}$ via labels $y_{1},\ldots,y_{n}$. It is clear that $(x_{1},y_{1}),\ldots,(x_{n},y_{n})$ are shattered in $\mathcal{C}_{\mathcal{Y}}$.
Suppose $\mathcal{C}_{\mathcal{Y}}$ has a binary compression scheme ($\kappa_{b}$,$\rho_{b}$), then we construct a compression scheme $(\kappa,\rho)$ for $\mathcal{C}$ as follows. Compression: Given a dataset $S=\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$ realizable by $\mathcal{C}$, construct $\kappa(S)$ as follows. Inflate $S$ to $\mathcal{S_{\mathcal{Y}}}$. Note that $S_{\mathcal{Y}}$ is realizable by $\mathcal{C}_{\mathcal{Y}}$, and $d_{\mathrm{VC}}=d_{\mathrm{G}}$. We can apply $\kappa_{b}$ to $S_{\mathcal{Y}}$ to get a compression of size $f(d_{\mathrm{G}})$. The compression points will be of the form $((x,y),z),z\in\{0,1\}$. For the points where $z=1$, we have that $(x,z)=(x_{i},y_{i})$ for some $i$, so we can add that to $\kappa(S)$, contributing $1$ for each. For the points where $z=0$, we need $\log|\mathcal{Y}|$ bits to add that to $\kappa(S)$. Thus, our compression size will be $\leq f(d_{\mathrm{G}})+f(d_{\mathrm{G}})\log|\mathcal{Y}|=O(f(d_{\mathrm{G}})\log|\mathcal{Y}|)$. Reconstruction: Our compression has enough information for us to retrieve the result of $\kappa_{b}(\mathcal{C}_{\mathcal{Y}})$. We can directly apply $\rho_{b}$ on this to get the desired result.
@@
Suppose all binary concept classes with VC dimension $d_{\mathrm{VC}}$ have a sample compression scheme of size $f(d_{\mathrm{VC}})$. Does every multiclass class with graph dimension $d_{\mathrm{G}}$ have a sample compression scheme of size $O\mathopen{}\left(f(d_{\mathrm{G}})\right)$?